# Understanding Token Probability Encoding in Output Embeddings

### v1

@@ 3.1 Algorithm
Based on the fact that the output probabilities are log-linearly encoded in a direction on the output embedding, and the correlation strengths significantly vary among dimensions, we propose an output probability editing algorithm as shown in Algorithm 1. We first calculate an output probability distribution $\mathbf{\alpha}_{\mathcal{D},\theta}$ on a corpus $\mathcal{D}$ and conduct MLR to calculate the $A_{\mathcal{D}}$ and the confidence ($p$-value) $\mathcal{S}$ of each element of the common encoding. Then we assign an editing weight $\Omega$ to every dimension in the embedding vector based on the significance of its correlation to the output probabilities, as shown in line 3 of Algorithm 1. We allocate more editing amount to stronger correlations to obtain smaller and more accurate editing [^2] Parameter $b$ is introduced to control the softness of such allocation, while the algorithm is stable on the parameter as shown in Appendix C. Basically, we suggest that a large $b$ is suitable for a large model.. Given the token index to be edited and the expected editing scale, we calculate the detailed editing amount on each dimension and update it as shown in line 4 of Algorithm 1.
As calculation costs, such an editing algorithm only needs a detect set $\mathcal{D}$ and a feed-forward process to calculate the $A_{\mathcal{D}}$ and $\mathcal{S}$. We are about to prove that it is stable for the size of $\mathcal{D}$ and consistently precise.
Table 2: Main results on the evaluation of Algorithm 1. Unedited: the baseline without any editing. Random: the baseline with a random $A_{\mathcal{D}}$. Shuffled: the baseline with a shuffled $A_{\mathcal{D}}$ from the origin one. It is difficult to conduct MAUVE experiments of GPT-J on such a repeating scale due to the enormous computational costs. $b$: the editing amount allocation parameter in Algorithm 1. Fine-grained results are presented in the Appendix D.
@@
6B -- 5 ${0.31}_{0.43}$ ${0.25}_{0.39}$ ${0.10}_{0.12}$ ${3.64}_{14.65}$-
@@
Figure 4: The expected probability scales against the actually edited scales measured in the edited LMs.
3.2 Experiment Settings
We use the same detect dataset $\mathcal{D}$ as in §2, and a set of scales of {1, 1.1, 1.2, 1.5, 2, 5, 10, 20} for both scaling up and down. We randomly select 10 tokens to be edited and conduct experiments on GPT2, GPT2-XL, and GPT-J.
Metrics.
We use 3 metrics to test the algorithm.
  • Scale error $e$: To describe the precision of the probability editing, given the expected editing scale $r$ and the actual measured edited scale $\hat{r}$ on a test dataset, the scale error is calculated as $e=|\ln(r)-\ln(\hat{r})|$.
  • KL divergence on the unedited token $d_{KL}^{r}$: To investigate the side effect on the unedited tokens, we calculate KL divergence between the probability distribution before and after editing on the unedited tokens.
  • MAUVE: To investigate the side effect on text generation, we generate a set of sentences from the edited model, then calculate the MAUVE [^3] Proposed by Pillutla et al. (2021), a measurement of the similarity of two language datasets. The value range is $[0,1]$, the larger means a greater similarity. with the generated set from the unedited model (see Appendix B.3).
Evaluations.
Following the widely-used aspects of model editing evaluations Yao et al. (2023), we define our evaluations as:
@@ 3.2 Experiment Settings
Figure 6: The MAUVE()($\uparrow$)( ↑ ) and KL divergence()($\downarrow$)( ↓ ) of output token probability distribution $d_{KL}$ against original distributions w.r.t. the dimension removing ratio on the output embedding. Solid curves: results of removing dimensions from the least important ones to the most important ones; Dashed curves: adversarial controlling experiment, removing dimensions reversely.
  • Reliability: The local effectiveness of the editing method. We use $e_{local}$, the scale error on the detect dataset $\mathcal{D}$.
  • Generalization: The global effectiveness of the editing method. We set multi-level generalization evaluations to ensure that the conclusions we obtain are the generalizable essence of the model: (i). the scale error on the other 8192 samples of WikiDpr as the In-domain Scale Error $e_{id}$; and (ii). the scale error on 2048 samples of BookCorpus Zhu et al. (2015) as the Out-of-domain Scale Error $e_{ood}$.
  • Specificity: The non-harmfulness to unedited part. We use two settings for this evaluation: (i). the averaged $d_{KL}^{r}$ on the three data samples (detect set, in-domain, out-of-domain), and (ii). the aforementioned MAUVE.
3.3 Results
The main statistical evaluation results are shown in Table 2, where we can get a basic intuition that the Algorithm 1 is accurate and generalizable, even on the out-of-domain data. Especially, despite the embedding tying in GPT2 and GPT2-XL, such an editing algorithm is still accurate and harmless, which demonstrates that the log-linear probability encoding is orthogonal to the possible semantical encoding in the output and also input embedding.
@@ 3.3 Results
Wide-scale stable: Large-scaled probability editing is supported by a global log-linear pattern.
The correlation of actually edited scales of token probability against the expected scales is shown in Fig. 4. The editing remains accurate even on a large scale of up to 20x. This indicates that the algorithm and the log-linear encoding are wide-ranging, not only locally effective, i.e. the log-linear encoding is a widely stable common essence.
@@
Instead of using the 8192 examples for the detecting dataset $\mathcal{D}$, we try various numbers of examples to test the generalization of the encoding on GPT2. As shown in Fig. 5, we find that even 2 examples can produce a distinct averaged probability distribution for precise probability editing. This result strengthens the significance and generalization of our findings, that is, effective statistical patterns can be found in a small set of samples.
These results demonstrate the wide-scale stability and generalization of our algorithm, while also reflecting the same attributes of token probability encoding in output embedding. We can confirm that such log-linear encoding of token probabilities is an inherent attribute of output embedding, which (1). has wide-scale linearity, allowing a large-scaling probability editing, and (2). is common among tokens, so only a small number of samples are needed to find an accurate encoding. By our editing experiments, the properties of probability encoding are strengthened.
@@ 4 Removing Dimensions with Weak Probability Encoding
Moreover, based on the sparsity of probability encoding as shown in Fig. 3, we can infer that a large number of dimensions are less effective for causal language modeling, where the expected role of the LM head is basically only to predict probabilities. So we try to reduce the less related dimensions towards a lightweight output head.
@@ 4.1 Method & Experiment Settings
First, we assign a weight of importance (or, saliency) to each dimension of output embedding from the aforementioned absolute MLR slopes as shown in (e-h) of Fig. 3. Then, we remove the dimensions in ascending order of such weight, i.e. we remove the less important dimension early. In detail, as a prototype setting in the laboratory, we only zero out the dimensions in the embedding matrix, while the dimensions of the attention mapping matrix, the feed-forward layer, and the dimensions of the last hidden state corresponding to the removed dimensions can also be removed equally.
@@
We use the same MLR settings as §2 on GPT2, Pythia 2.8B, and GPT-J. As metrics, we test the MAUVE of the pruned model similarly to §3.2, and test the KL divergence of the averaged probability distribution against the original model on the out-of-domain dataset mentioned before. As an adversarial controlling experiment, we remove dimensions in descending order of importance, i.e. we remove the more important dimension early, reverse to the original settings.
@@ 4.2 Result
The results are shown in Fig. 6. With the ascending removing order (the positive experiment, solid curve), at the beginning of the removal, both metrics deteriorate slightly until around 50% of the dimensions are removed. Regarding $\mathrm{MAUVE}>0.80$ as a threshold, we can confirm that $30\%\sim 40\%$ of the dimensions in the output embedding can be removed without significant impairment of the causal language modeling ability of LMs. In contrast, in the adversarial experiment with a descending removing style (dashed curve), both metrics deteriorate sharply at the beginning of removing. Such results suggest that the weights from the MLR are faithful importance metrics.
Noticeably, our method also works on the tied models, i.e. removing unnecessary dimensions concurrently from the input embedding and the output embedding, the causal language modeling ability of the LM remains at a considerable level.
Consistent with previous works Kovaleva et al. (2021); Timkey and van Schijndel (2021); Gordon et al. (2020), our results identify the difference in the importance among the dimensions of the output embedding. Moreover, we confirm that the MLR results of the log-linear correlation are a good metric of importance, i.e. saliency score. This can be a new paradigm of saliency if a direct one-step statistical model can be found between the features and the model output.
@@ 5 Output Embedding Learns Corpus Frequency in Early Training Steps
Additionally, the findings in this paper inspire us to utilize such log-linear correlation for detecting the encoding of corpus token frequency in the output embedding during the training. Intuitively, since the model can refract overall output probability distribution in the output embedding matrix, it should be forced to produce the same output distribution with the corpus token frequency by the training objective. Therefore, when a log-linear encoding of token frequency of the training corpus is observed in the output embedding, we can confirm that the output embedding learns the token frequency information of the corpus.
We use the Pythia suite Biderman et al. (2023), where sequences of the model intermediate training checkpoints are saved. We estimate the token frequencies of the training corpus Pile Gao et al. (2020); Biderman et al. (2022) by sampling 14.3B tokens, then conduct MLR on the negative logarithm of the token frequencies w.r.t. the output embeddings on each training checkpoint. The results are shown in the left part of Fig. 7, where we can confirm an effective encoding since the very early steps of the training process, and larger models have slightly better fitting goodness but almost no difference in the timing of emergence.
Moreover, we want to know whether such a phenomenon is a subsidiary effect of the convergence of the parameters. We use the convergence rate following the Chen et al. (2022) to describe the actual training completion: denoting the trained parameter matrix as $\theta^{*}$, the initialized parameter matrix as $\theta_{0}$, and the parameter matrix at step $t$ as $\theta_{t}$, the convergence rate at step $t$ can be written as $(1-\|\theta^{*}-\theta_{t}\|_{F}/\|\theta^{*}-\theta_{0}\|_{F})$. We visualize the convergence rate of the input and output embeddings and the query-key-value mapping matrix of the multi-head attention blocks on Pythia-1B, as shown in the right part of Fig. 7. We find that the log-linear correlation occurs even earlier than an obvious convergence trend is observed by the convergence rate. Or rather, the appearance of such correlation happens to be the starting point of the convergence process of the model. We infer that the output embedding should learn the coarse-grained output pattern earlier than the semantics details.
@@ Conclusion.
In this paper, based on the observation of a linear-like correlation between the output token probability and the output token embeddings shown in Fig. 1, we derive an approximate log-linear correlation from the nature of $\mathrm{softmax}$ output head with a large output space and concentrated output value, i.e. the output token probabilities are encoded in a common direction of output embeddings. Along such encoding direction, we edit the token probabilities with high accuracy, stability, and generalizability. Then, based on the sparsity of the encoding, we can distinguish the contributions of dimensions for the output probability of the model. We try removing the determined non-contributing dimensions, and no critical deterioration is found. Finally, based on the findings, we find that the LMs catch the token frequency in training data at very early steps in the training process log-linearly, even earlier than an obvious convergence trend is observed. This paper reveals the inner mechanism of LM heads on the causal language modeling task and helps understand the global principles and training dynamics of LMs.
@@
Previous work about analyzing LM heads was conducted by Kobayashi et al. (2023), where a correlation between the output token probability and bias term in the LM head was found. They declared that the bias term is a projection to extract the probability from the output embedding, but no more discussion about the embedding matrix, which is the major component of the LM head. Making up for their work is one of the original motivations for our work. More discussions about related works can be found in Appendix A.
@@
Our output probabilities editing algorithm on the LM head reveals the possibility of model editing on only the output probability rather than in the hidden states or the lower layers Dai et al. (2022); Meng et al. (2022), which is more concise, and easy to explain. Some global toxicity generation Gehman et al. (2020), or biases in some application scenario Fei et al. (2023) can be suppressed by such a method, but as we will mention in the Limitations, it is elegant but not engineering-oriented. Moreover, the dimension compression method in this paper can be an easy-to-use and harmless inference-time acceleration. Notice that we have a supervising on such dimension compression by the MLR slopes, so it can be more accurate than the previous random or unsupervised pruning Gordon et al. (2020).
@@
Saliency score Zhao et al. (2024); Sun et al. (2021) is to weight a component (feature or parameter) in a deep learning model by its importance. In this paper, we find that the log-linear token probability encoding works like a saliency score towards the output embedding. We build a log-MLR model to assign saliency scores to the parameters, and such a statistical method can be a new paradigm of model-based saliency Dabkowski and Gal (2017), if strong one-step correlations can be found between the output and components of the neural network, a closed-form saliency model can be proposed instead of a universal statistical model.
Duality between input and output embeddings.
In addition, we find some duality between input and output embeddings, e.g., they both have a good log-linearity w.r.t. the output probability and almost overlapping training curves. Further work can be focused on such a phenomenon to get a better understanding of the inner states, and the interaction between components of LMs.

### v2

@@ 3.1 Algorithm
Table 2: Main results on the evaluation of Algorithm 1. Unedited: the baseline without any steering. Random: the baseline with a random $A_{\mathcal{D}}$. Shuffled: the baseline with a shuffled $A_{\mathcal{D}}$ from the calculated value. $b$: the softness parameter in Algorithm 1. Fine-grained results are reported in the Appendix C.
@@
6B 5 ${0.31}_{0.43}$ ${0.25}_{0.39}$ ${0.10}_{0.12}$ ${3.64}_{14.65}$- [^3] It is difficult to conduct MAUVE experiments of GPT-J on such a repeating scale due to computational costs.
Based on the fact that the output probabilities are log-linearly encoded in a direction on the output embedding sparsely, we propose an output probability steering algorithm as shown in Algorithm 1. Step 1: We first estimate an averaged output probability distribution $\mathbf{\alpha}_{\mathcal{D},\theta}$ on a detect dataset $\mathcal{D}$ and conduct MLR to calculate the encoding direction $A_{\mathcal{D}}$ and the correlation significance ($p$-value used) $\mathcal{S}$ of each element of the common encoding. Step 2: Then we assign a steering weight $\Omega$ to every dimension in the embedding vector based on the significance of its correlation to the output probabilities, as shown in line 3 of Algorithm 1. We allocate more steering amount to stronger correlations to obtain smaller and more accurate steering [^4] Parameter $b$ is introduced to control the softness, or sparsity of such allocation, while the algorithm is stable on the parameter as shown in Appendix B. Basically, we suggest that a large $b$ is suitable for a large model.. Step 3: Given the token index to be steered and the expected steering scale, we calculate the detailed steering amount on each dimension and update it as shown in line 4 of Algorithm 1.
As data costs, our steering only needs a detect set $\mathcal{D}$ and a feed-forward process to calculate the $A_{\mathcal{D}}$ and $\mathcal{S}$, and we are about to prove that it is stable for the size of $\mathcal{D}$ and consistently precise.
3.2 Experiment Settings
We use the same detect dataset $\mathcal{D}$ as in §2, and a set of scales of {1, 1.1, 1.2, 1.5, 2, 5, 10, 20} for both scaling up and down. We randomly select 10 tokens to be steered and conduct experiments on GPT2, GPT2-XL, and GPT-J.
Metrics.
We use 3 metrics to test the algorithm.
  • Scale error $e$: To describe the precision of the probability steering, given the expected steering scale $r$ and the actual measured steered scale $\hat{r}$ on a test dataset, the scale error is calculated as $e=|\log(r)-\log(\hat{r})|$.
  • KL divergence on the retained token $d_{\mathrm{KL}}^{r}$: To investigate the side effect on the retained tokens, we calculate KL divergence between the probability distribution before and after steering with the steered token excluded.
  • MAUVE: To investigate the side effect on text generation, we generate a set of sentences from the steered model, then calculate the MAUVE [^5] Proposed by Pillutla et al. (2021), a measurement of the similarity of two language datasets. The value range is $[0,1]$, the larger means a greater similarity. with the generated set from the original model (see Appendix A.3).
Evaluations.
Applying the aforementioned metrics, following the widely-used aspects of model editing evaluations Yao et al. (2023), we define our evaluations in 3 categories:
  • Reliability: The local effectiveness of the steering. We use $e_{\mathrm{local}}$, the scale error with the detect dataset $\mathcal{D}$ as the test dataset.
  • Generalization: The global effectiveness of the steering. We set multi-level generalization evaluations to ensure robustness: (i) In-domain Scale Error $e_{\mathrm{id}}$: the scale error with the other 8192 samples of WikiDpr (where the updated parameters are estimated) as the test dataset; and (ii) Out-of-domain Scale Error $e_{\mathrm{ood}}$: the scale error on 2048 samples of BookCorpus Zhu et al. (2015) as the test dataset.
  • Specificity: The harmless to the retained part. We use two metrics for this evaluation: (i) the averaged $d_{\mathrm{KL}}^{r}$ on the three data samples (detect set, in-domain, out-of-domain), and (ii) the aforementioned MAUVE.
3.3 Results
The main statistical evaluation results are shown in Table 2. Compared to the baseline values, the results from Algorithm 1 are more accurate, generalizable, and harmless, even on the out-of-domain data. Especially, despite the embedding tying in GPT2 and GPT2-XL, such a steering algorithm is still accurate and harmless, which demonstrates that the log-linear probability encoding is orthogonal to the possible semantical encoding in the output and also input embedding.
@@ 3.3 Results
Figure 5: The expected probability scales against the actually steered scales measured in the steered LMs.
@@
Figure 6: The MAUVE()($\uparrow$)( ↑ ) and KL divergence()($\downarrow$)( ↓ ) of output token probability distribution $d_{\mathrm{KL}}$ against original distributions w.r.t. the dimension removing ratio on the output embedding. Solid curves: results of removing dimensions from the least important ones to the most important ones; Dashed curves: adversarial controlling experiment, removing dimensions reversely.
@@
Wide-scale stable: Large-scaled probability steering is supported by a global log-linear pattern.
The correlation of actually steered scales of token probability against the expected scales is shown in Fig. 5. The steering remains accurate even on a large scale of up to 20x. This indicates that the algorithm and the log-linear encoding are wide-ranged, not only locally effective, i.e. the log-linear encoding is a widely stable common essence.
@@
Instead of using the 8192 examples for the detecting dataset $\mathcal{D}$, we try various numbers of examples to test the generalization of the encoding on GPT2. As shown in Fig. 4 and 16, we find that even 2 examples can produce a distinct averaged probability distribution for precise probability steering. This result strengthens the significance and generalization of our findings, that is, effective encodings can be detected in a small set of samples.
These steering results demonstrate a clear causality between the log-linear encoding and the token probabilities, further than a simple statistical covariance relationship, and also demonstrate a wide-scale stability and generalization of our algorithm, while reflecting the same attributes of token probability encoding in output embedding. We can confirm that such log-linear encoding of token probabilities is an inherent attribute of output embedding, which (1) has wide-scale linearity, allowing a large-scaling probability steering, and (2) is common among tokens, so only a small number of samples are needed to find an accurate encoding. By our steering experiments, the properties of probability encoding are strengthened.
@@ 4 Removing Dimensions with Weak Probability Encoding
Moreover, based on the sparsity of probability encoding as shown in Fig. 2, we can infer that a large number of dimensions in the output embeddings are less effective for next token decoding. So we try to reduce the less related dimensions towards a lightweight output head.
@@ 4.1 Method & Experiment Settings
First, we assign the aforementioned absolute MLR slopes as shown in (e-h) of Fig. 2 as the importance (or, saliency) to each dimension of output embedding. Then, we remove the dimensions in ascending order of such weight, i.e. we remove the identified less important dimension early. In detail, as a prototype setting in the laboratory, we only zero out the dimensions in the embedding matrix, while the dimensions of the attention mapping matrix, the feed-forward layer, and the dimensions of the last hidden state corresponding to the removed dimensions can also be removed equally.
@@
We use the same MLR settings as §2 on GPT2, Pythia 2.8B, and GPT-J. As metrics for the causal language modeling quality, we test the MAUVE of the pruned model similarly to §3.2, and test the KL divergence of the averaged probability distribution against the original model on the out-of-domain dataset mentioned before. As an adversarial controlling experiment, we remove dimensions in descending order of importance, i.e. we remove the more important dimension early, reverse to the original settings.
@@ 4.2 Result
The results are shown in Fig. 6. With the ascending removing order (the positive experiment, solid curve), at the beginning of the removal, both metrics deteriorate slightly until around 50% of the dimensions are removed. Regarding $\mathrm{MAUVE}>0.80$ as a threshold, we can confirm that $30\%\sim 40\%$ of the dimensions in the output embedding can be removed without significant impairment of the causal language modeling quality of LMs. In contrast, in the adversarial experiment with a descending removing style (dashed curve), both metrics deteriorate sharply at the beginning of removing. Such results suggest that the weights from the MLR are faithful importance metrics.
Noticeably, our method also works on the tied models (Fig. 6 (a)), i.e. removing unnecessary dimensions concurrently from the input embedding and the output embedding, the metrics remain at a considerable level.
Consistent with previous works Kovaleva et al. (2021); Timkey and van Schijndel (2021); Gordon et al. (2020), our results identify the anisotropy in the importance among the dimensions of the output embedding. Moreover, we confirm by dimensional ablation that the MLR results of the log-linear correlation are a good metric of importance, i.e. saliency score. This can be a new paradigm of saliency if a direct one-step statistical model (like MLR in our paper) can be estimated between the features and the model output.
@@ 5 Output Embedding Learns Corpus Frequency in Early Training Steps
Additionally, the findings in this paper inspire us to utilize such log-linear correlation for detecting the encoding of corpus token frequency in the output embedding during the pre-training of LMs, for a closer observation of the pre-training dynamics. Intuitively, since the model can refract overall output probability distribution in the output embedding matrix, it should be forced to produce the same output distribution with the corpus token frequency by the training objective. Therefore, when a log-linear encoding of token frequency of the training corpus is observed in the output embedding, we can confirm that the output embedding learns the token frequency information of the corpus.
We use the Pythia suite Biderman et al. (2023), where sequences of the model intermediate training checkpoints are accessible. We estimate the token frequencies of the pre-training corpus Pile Gao et al. (2020); Biderman et al. (2022) by sampling 14.3B tokens, then conduct MLR on the negative logarithm of the estimated token frequencies w.r.t. the output embeddings on each training checkpoint, with various training steps. The results of MLR fitting goodness against training steps are shown in the left part of Fig. 7, where we can confirm an effective encoding since the very early steps of the training process, and larger models have slightly better fitting goodness but almost no difference in the timing of emergence.
Moreover, we want to know whether such a phenomenon is a subsidiary or subsequent effect of the convergence of the parameters. We use the convergence rate following the Chen et al. (2022) to describe the actual training completion: denoting the trained parameter matrix as $\theta^{*}$, the initialized parameter matrix as $\theta_{0}$, and the parameter matrix at step $t$ as $\theta_{t}$, the convergence rate at step $t$ can be written as $(1-\|\theta^{*}-\theta_{t}\|_{F}/\|\theta^{*}-\theta_{0}\|_{F})$. We visualize the convergence rate of the input and output embeddings and the query-key-value mapping matrix of the multi-head attention blocks on Pythia-1B, as shown in the right part of Fig. 7. We find that the log-linear correlation occurs even earlier than an obvious convergence trend is observed by the convergence rate, and the appearance of such correlation happens to be the starting point of the convergence process of the model. We infer that the output embedding should learn the coarse-grained output pattern earlier than the semantics details.
@@ Conclusion.
In this paper, based on the observation of a linear-like correlation between the output token probability and the output token embeddings shown in Fig. 1, we derive an approximate log-linear correlation from the nature of $\mathrm{softmax}$-based output head with a large output space and concentrated output value, i.e. the output token probabilities are encoded in a common direction of output embeddings. To test the causality of such correlation, we steer the token probabilities along such encoding direction with high accuracy, stability, and generalizability. Then, based on the sparsity of the encoding, we can distinguish the contributions of dimensions for the output probability of the model, which allows us to remove the determined non-contributing dimensions, and no critical deterioration is found. Finally, based on the findings, we find that the LMs catch the token frequency in training data at very early steps in the training process log-linearly, even earlier than an obvious convergence trend is observed. This paper reveals the inner mechanism of LM heads on the causal language modeling task and helps understand the global principles and training dynamics of LMs.
@@
Previous work about analyzing LM heads was conducted by Kobayashi et al. (2023), where a correlation between the output token probability and bias term in the LM head was found. They declared that the bias term is a projection to extract the probability from the output embedding, but almost no more discussion about the embedding matrix, which is the major component of the LM head. Making up for their work is one of the original motivations for our work. More discussions about related works can be found in §7.
@@
Our output probabilities steering algorithm on the LM head reveals the possibility of model steering on only the output probability rather than in the hidden states or the lower layers Dai et al. (2022); Meng et al. (2022), which is more concise, and easy to explain. Some global toxicity generation Gehman et al. (2020), or biases in some application scenario Fei et al. (2023) can be suppressed by such a method, but as we will mention in the Limitations, it is elegant but not engineering-oriented. Moreover, the dimension compression method in this paper can be an easy-to-use and harmless inference-time acceleration. Notice that we have supervision on such dimension compression by the MLR slopes, so it can be more accurate than the previous random or unsupervised pruning Gordon et al. (2020).
@@
Saliency score Zhao et al. (2024); Sun et al. (2021) is to weight a component (feature or parameter) in a deep learning model by its contribution. In this paper, we find that the log-linear token probability encoding works like a saliency score towards the output embedding, and build a log-MLR model to assign saliency scores to the parameters. Such a posterior statistical method can be a new paradigm of model-based saliency Dabkowski and Gal (2017), if one-step correlations can be detected between the output and components of the neural network, a closed-form saliency model can be proposed instead of a universal statistical model.
7 Related Works
Geometry of Input Embedding.
As a similar research object with the output embedding, it was found that the word embeddings in LMs, as well as the hidden states, are anisotropy Mu and Viswanath (2018); Ethayarajh (2019); Gao et al. (2018), i.e., these vectors share a common radial bias. Such anisotropies hurt the expressiveness of word embeddings, and the word frequency in the corpus may be an inducement Mu and Viswanath (2018); Valentini et al. (2023). Also, some efforts tried to remove the harmfulness of anisotropies and towards isotropy word embeddings Mu and Viswanath (2018); Gong et al. (2018). These works are based on input embedding, while our paper is on output embedding. Although we can confirm that the input and output embeddings act similarly, they are still completely different components of untied LMs. So the existing conclusions on input embeddings cannot overwrite our work.
Embedding Tying in LMs.
(Language) Model Editing and Model Pruning.
Training Dynamics (of LMs).
Investigating what is happening in the training process of language models and other deep learning models is an attractive research topic. Many works about training trajectory Kalra and Barkeshli (2024); Jastrzębski et al. (2018); Lewkowycz et al. (2020), early period training behaviors Frankle et al. (2019); Kalra and Barkeshli (2024); Achille et al. (2018), loss landscape Neyshabur et al. (2020); Keskar et al. (2017); Li et al. (2018), and “knowledge” earned in different stages of training Tirumala et al. (2022); Liu et al. (2021) have been done. Differences in training speed among various network layers (the deeper-slower pattern) have been discovered by Chen et al. (2022), while in this paper, following their method, we don’t find a similar pattern.